# TikTok’s recommendations skewed towards Republican content during the 2024 U.S. presidential race

### v1

@@ 10.1. None of the above topics
Verified channel -0.5313 0.5879 0.5029 0.6871 p <0.001
Follower count 0.2907 1.3374 1.2364 1.4466 p <0.001
Video count 0.0473 1.0484 0.9835 1.1176 p = 0.1469
Recommendation video metrics
Comment count 0.0795 1.0828 1.0071 1.1642 p = 0.0316
Share count 0.0529 1.0543 0.971 1.1447 p = 0.2081
Likes per recommendation -0.4145 0.6607 0.6087 0.7171 p <0.001
Comments per recommendation 0.1954 1.2158 1.1037 1.3392 p <0.001
Conditioning metrics
Engagement score of conditioning videos -0.1975 0.8208 0.7459 0.9032 p <0.001
Number of conditioning videos -0.1148 0.8915 0.8374 0.9491 p <0.001
Table S23: Logistic regression coefficients and odds-ratios predicting whether a given video watch is mismatched/cross-party aligned.
@@
Georgia 0.0526 1.054 0.8296 1.3392 p = 0.6668
Texas 0.2132 1.2376 0.9662 1.5854 p = 0.0915
@@
Week of experiment -0.0361 0.9645 0.9478 0.9815 p <0.001
@@
Video count -0.2109 0.8098 0.6876 0.9538 p = 0.0115
Follower count 0.251 1.2854 1.1552 1.4302 p <0.001
Cumulative like count 0.2126 1.2369 1.0837 1.4116 p = 0.0016
Recommendation video metrics
Likes per recommendation -0.34 0.7117 0.6276 0.8072 p <0.001
Comments per recommendation 0.1591 1.1725 1.0489 1.3106 p = 0.0051
Conditioning metrics
Engagement score of conditioning videos -0.2465 0.7815 0.6706 0.9108 p = 0.0016
Number of conditioning videos -0.036 0.9646 0.8635 1.0776 p = 0.5240
Table S24: Logistic regression coefficients and odds-ratios predicting whether a given video watch is mismatched/cross-party aligned when only considering videos published by top Republican and Democrat channels.
Variable Coefficient Odds Ratio OR Lower CI OR Upper CI P-value
Bot conditioning (ref = Rep.)
Democrat 1.0611 2.8896 2.3823 3.5049 p <0.001
Partisanship type (ref = Positive)
Negative partisanship 0.5788 1.7838 1.5433 2.0619 p <0.001
State (ref = NY)
Georgia -0.0352 0.9655 0.8325 1.1197 p = 0.6420
Texas 0.0692 1.0717 0.9237 1.2433 p = 0.3612
Time
Week of experiment -0.0444 0.9566 0.9478 0.9655 p <0.001
Recommendation channel metrics
Verified channel -0.5176 0.5959 0.5096 0.6969 p <0.001
Video count 0.0141 1.0142 0.9492 1.0836 p = 0.6771
Follower count 0.3346 1.3973 1.2872 1.5169 p <0.001
Recommendation video metrics
Comment count 0.0662 1.0684 0.9922 1.1505 p = 0.0796
Share count 0.045 1.046 0.9597 1.14 p = 0.3056
Likes per recommendation -0.4191 0.6577 0.6054 0.7144 p <0.001
Comments per recommendation 0.2218 1.2483 1.1314 1.3773 p <0.001
Conditioning metrics
Engagement score of conditioning videos -0.1982 0.8202 0.7456 0.9024 p <0.001
Number of conditioning videos -0.1203 0.8866 0.8325 0.9443 p <0.001
Table S25: Logistic regression coefficients and odds-ratios predicting whether a given video watch is mismatched/cross-party aligned when accounting for the partisanship of the video.
@@
Table S26: Engagement metrics of videos watched during recommendation stage.
@@
Table S27: Engagement metrics of channels watched during recommendation stage.
@@
Table S28: Engagement metrics of videos watched during conditioning stage.
@@
Table S29: Engagement metrics of channels watched during conditioning stage.
Additional Supplementary Figures
Figure S2: A device’s timeline during a weekly experimental run
@@
Figure S3: Rolling average of political content viewed by bots of different conditioning in 10 video windows.
@@
Figure S4: Rolling average of political content about the US elections or about major political candidates (out of all political content) viewed by bots of different conditioning in 10 video windows.
@@
Figure S5: Rolling average of political content that is ideologically matched (out of all political content) to bots of different conditioning in 10 video windows. Pro Democrat or Anti Republican videos are ideologically matched with Democrat bots, while Pro Republican or Anti Democrat videos are ideologically matched with Republican bots.
@@
Figure S6: Rolling average of political content that is ideologically mismatched (out of all political content) to bots of different conditioning in 10 video windows. Pro Republican or Anti Democrat videos are ideologically mismatched with Democrat bots, while Pro Democrat or Anti Republican videos are ideologically mismatched with Republican bots.
@@
Figure S7: Of topic-stance pairs with at least 100 videos, the difference in the proportion of videos on the topic with the given stance and the proportion of videos of that stance out of all political videos. Positive values indicate a higher than average representation of the stance within a topic, while negative values lower than average representation of the stance within the topic. Plots A, B, C, D, E correspond to the Pro Democrat, Anti Republican, Anti Democrat, Pro Republican, and Neutral stances, respectively.
@@
Figure S8: (A, B) The proportion of videos on a given topic which are ideologically-aligned, ideologically-opposing, or neutral, seen by Democrat- and Republican-conditioned bots, respectively. For each plot, topics are listed in descending order of ideological-alignment.

### v2

@@ Acknowledgments
    Verified channel -0.2995 0.7412 0.5673 0.9684 p = 0.0281
    Video count 0.0768 1.0798 0.9595 1.2152 p = 0.2027
    Follower count 0.1158 1.1228 0.9892 1.2745 p = 0.0732
    Supplementary Table 14: Logistic regression coefficients and odds-ratios predicting whether a given video watch is mismatched/cross-party aligned when only considering videos published by top Republican and Democrat channels.
@@
    Georgia -0.011 0.9891 0.7844 1.2471 p = 0.9259
    Texas 0.0664 1.0687 0.8464 1.3494 p = 0.5767
    Partisanship x Bot conditioning
    Negative partisanship × Republican bot 0.2881 1.3339 0.9091 1.9573 p = 0.1408
    Negative partisanship × Democrat bot 1.3051 3.6881 2.6652 5.1037 p <0.001
    Top account x Bot conditioning
    Top account × Republican bot 0.0433 1.0442 0.7113 1.5331 p = 0.8251
    Top account × Democrat bot -1.0339 0.3556 0.2651 0.4771 p <0.001
@@
    Week of experiment -0.0336 0.967 0.9506 0.9836 p <0.001
    Conditioning metrics
    Engagement score of conditioning videos -0.1176 0.889 0.7566 1.0446 p = 0.1529
    Number of conditioning videos -0.0721 0.9304 0.8384 1.0325 p = 0.1746
    Recommendation video metrics
    Likes per recommendation -0.2262 0.7975 0.7043 0.903 p <0.001
    Video length -0.2116 0.8093 0.5988 1.0938 p = 0.1686
    Share count -0.1223 0.8849 0.7914 0.9894 p = 0.0317
    Comments per recommendation 0.1199 1.1274 0.9562 1.3293 p = 0.1537
    Opposing comment proportion 0.8272 2.287 2.0563 2.5436 p <0.001
@@
    Verified channel -0.2846 0.7523 0.575 0.9843 p = 0.038
    Video count 0.0449 1.046 0.9261 1.1814 p = 0.4693
    Follower count 0.1711 1.1866 1.0401 1.3538 p = 0.0109
    Supplementary Table 15: Logistic regression coefficients and odds-ratios predicting whether a given video watch is mismatched/cross-party aligned when accounting for the partisanship of the video.
@@
    Supplementary Table 16: Engagement metrics of videos watched during recommendation stage.
@@
    Supplementary Table 17: Engagement metrics of channels watched during recommendation stage.
@@
    Supplementary Table 18: Engagement metrics of videos watched during conditioning stage.
@@
    Supplementary Table 19: Engagement metrics of channels watched during conditioning stage.
    Free response questions
    Variable
    Mentioned changes
    to political content
    when asked about
    content overall
    Mentioned shifts
    towards co-partisan
    content when asked
    about content overall
    Mentioned shifts
    towards co-partisan
    content when asked
    about political content
    Mentioned shifts to
    positive content
    Intercept -1.286** 0.011 -0.094 -0.202
    Republican Party (ref. Democratic Party) -0.353* 0.024 0.08** 0.274***
    Age (ref. 65+ years old)
    18-24 years old 0.099 -0.04 0.071 -0.035
    25-34 years old 0.486 -0.027 0.036 0.051
    35-44 years old 0.464 -0.021 0.024 0.044
    45-54 years old 0.718 -0.003 -0.001 0.011
    55-64 years old 0.364 -0.015 0.062 -0.072
    Race (ref. White)
    Asian 1.013** 0.1 0.152 -0.266*
    Black -0.074 0.002 0.025 0.001
    Hispanic 0.243 -0.03 0.06 -0.377**
    Native American -0.558 -0.005 -0.038 0.144
    Two or more races 0.891 -0.0 -0.089 -0.456*
    Other -0.123 -0.003 -0.009 0.242
    Prefer not to answer -237.16*** 0.009 0.016 0.15*
    Female (ref. Male) 0.036 0.007 0.042 0.0
    Education (ref. Bachelor’s degree)
    Associates or technical degree 0.381 0.009 0.07 -0.247*
    Graduate or professional degree -0.345* -0.006 0.026 0.078
    High school diploma or GED -0.38 -0.027 -0.071 -0.224*
    Some college, but no degree 0.133 -0.006 -0.017 -0.119
    Some high school or less -0.75 -0.013 0.175 0.367
    Prefer not to say -240.424*** -0.021 -0.052 0.883*
    Supplementary Table 20: Regression coefficients for various models estimating responses to free-response (text-entry) questions based on participant demographics; (Q1 - Q3) of Survey. * : $p<0.05$; **: $p<0.01$; ***: $p<0.001$
    Structured questions
    Variable
    More co-partisan/
    cross-partisan
    content
    More Pro-Trump
    More Anti-Trump
    content
    More political/
    Less political
    content
    More positive/
    More negative
    content
    More optimistic/
    More pessimistic
    content
    More content
    you agree
    /disagree with
    Intercept 0.394 4.755*** 7.25*** 4.68*** 4.575*** 6.267***
    Republican Party (ref. Democratic Party) 1.445*** 1.071*** -0.047 1.029*** 0.994*** 0.358*
    Age (ref. 65+ years old)
    18-24 years old 0.38 0.174 -0.13 -0.38 0.114 -0.018
    25-34 years old -0.236 -0.172 -0.388 -0.527 -0.091 -0.376
    35-44 years old -0.416 0.073 -0.29 -0.628 -0.026 -0.386
    45-54 years old -0.677 -0.166 -0.491 -0.698 -0.219 -0.895*
    55-64 years old -0.228 -0.639 -0.077 -0.531 -0.019 -0.368
    Race (ref. White)
    Asian 0.356 -1.156* -1.112* -1.012* -1.726*** -0.497
    Black 0.139 -0.329 0.164 0.062 0.263 0.02
    Hispanic -0.751 0.418 0.418 -0.516 0.018 -0.251
    Native American 0.715 -0.702 -0.204 0.619 1.253 0.043
    Two or more races -1.812* 0.317 0.476 -1.49* -1.123 -2.146**
    Other 0.583 1.428 -1.188 1.182 1.811 0.657
    Prefer not to answer -0.152 -1.416 2.507* -2.141 -2.587 -4.043
    Female (ref. Male) -0.006 -0.167 -0.369* -0.012 0.103 0.152
    Education (ref. Bachelor’s degree)
    Associates or technical degree 0.337 -0.555 -0.787 -0.325 -0.562 -0.274
    Graduate or professional degree 0.636*** 0.172 0.341* 0.424* 0.541** 0.546**
    High school diploma or GED -0.332 -0.765* -0.134 -0.73* -0.367 -0.246
    Some college, but no degree 0.362 -0.364 -0.663 -0.075 -0.414 -0.047
    Some high school or less 0.66 -0.969 -3.007* -0.079 -0.567 -0.725
    Prefer not to say 2.583*** 2.269*** -5.543* 3.932*** -4.645 2.609*
    Supplementary Table 21: Regression coefficients for various models estimating responses to structured (scale-based) questions based on participant demographics; (Q4 - Q9) of Survey. * : $p<0.05$; **: $p<0.01$; ***: $p<0.001$
    Demographic Count
    25-34 years old 373
    35-44 years old 232
    45-54 years old 173
    18-24 years old 109
    55-64 years old 86
    65+ years old 35
    Female 547
    Male 449
    Non-binary / third gender 12
    White 646
    Black 276
    Asian 28
    Hispanic 25
    Two or more races 21
    Native American 7
    Other 5
    Bachelor’s degree 417
    Graduate or professional degree 412
    Some college, but no degree 74
    High school diploma or GED 55
    Associates or technical degree 46
    Some high school or less 3
    Prefer not to say 1
    Party
    Republican Party 552
    Democratic Party 317
    Independent 139
    Democrat Strong
    Strong Democrat 210
    Not so strong Democrat 106
    Prefer not to answer 1
    Republican Strong
    Strong Republican 346
    Not so strong Republican 202
    Prefer not to answer 4
    Independent close to Democrat or Republican
    Neither 66
    Democratic Party 35
    Republican Party 34
    Not sure 3
    Prefer not to answer 1
    Liberal/Conservative
    Conservative 341
    Liberal 215
    Moderate 178
    Very conservative 160
    Very liberal 108
    Not sure 5
    Prefer not to answer 1
    Voted in 2024 elections
    Yes 924
    No 64
    Prefer not to answer 20
    Party in 2024 elections
    Republican 590
    Democrat 307
    Independent 15
    Other 6
    Libertarian 4
    Green 2
    Plans to vote in 2026 elections
    Yes 798
    Not sure 152
    No 46
    Prefer not to answer 12
    Party in 2026 elections
    Republican 449
    Democrat 290
    Independent 39
    Other 15
    Libertarian 4
    Green 1
    Supplementary Table 22: Participant demographic counts
    Supplementary Figures
    Supplementary Figure 1: A device’s timeline during a weekly experimental run
@@
    Supplementary Figure 2: Rolling average of political content viewed by bots of different conditioning in 10 video windows.
@@
    Supplementary Figure 3: Rolling average of political content about the US elections or about major political candidates (out of all political content) viewed by bots of different conditioning in 10 video windows.
@@
    Supplementary Figure 4: Rolling average of political content that is ideologically matched (out of all political content) to bots of different conditioning in 10 video windows. Pro Democrat or Anti Republican videos are ideologically matched with Democrat bots, while Pro Republican or Anti Democrat videos are ideologically matched with Republican bots.
@@
    Supplementary Figure 5: Rolling average of political content that is ideologically mismatched (out of all political content) to bots of different conditioning in 10 video windows. Pro Republican or Anti Democrat videos are ideologically mismatched with Democrat bots, while Pro Democrat or Anti Republican videos are ideologically mismatched with Republican bots.
@@
    Supplementary Figure 6: Of topic-stance pairs with at least 100 videos, the difference in the proportion of videos on the topic with the given stance and the proportion of videos of that stance out of all political videos. Positive values indicate a higher than average representation of the stance within a topic, while negative values lower than average representation of the stance within the topic. Plots A, B, C, D, E correspond to the Pro Democrat, Anti Republican, Anti Democrat, Pro Republican, and Neutral stances, respectively.
@@
    Supplementary Figure 7: (A, B) The proportion of videos on a given topic which are ideologically-aligned, ideologically-opposing, or neutral, seen by Democrat- and Republican-conditioned bots, respectively. For each plot, topics are listed in descending order of ideological-alignment.